### v5

# Global ${\mathcal{L}}^{2}$ minimization at uniform exponential rate via geometrically adapted gradient descent in Deep Learning

Abstract.
We consider the scenario of supervised learning in Deep Learning (DL) networks, and exploit the arbitrariness of choice in the Riemannian metric relative to which the gradient descent flow can be defined (a general fact of differential geometry). In the standard approach to DL, the gradient flow on the space of parameters (weights and biases) is defined with respect to the Euclidean metric. Here instead, we choose the gradient flow with respect to the Euclidean metric in the output layer of the DL network. This naturally induces two modified versions of the gradient descent flow in the parameter space, one adapted for the overparametrized setting, and the other for the underparametrized setting. In the overparametrized case, we prove that, provided that a rank condition holds, all orbits of the modified gradient descent drive the ${\mathcal{L}}^{2}$ cost to its global minimum at a uniform exponential convergence rate; one thereby obtains an a priori stopping time for any prescribed proximity to the global minimum. We point out relations of the latter to sub-Riemannian geometry. Moreover, we generalize the above framework to the situation in which the rank condition does not hold; in particular, we show that local equilibria can only exist if a rank loss occurs, and that generically, they are not isolated points, but elements of a critical submanifold of parameter space.
MSC classes: 57R70, 62M45
1. Introduction and Main Results
Use of the gradient descent flow generated by the cost (loss) function is ubiquitous as a method for cost minimization in the training of Deep Learning (DL) networks. While time proven and very successful in general, its implementation exhibits inherent challenges, including the trapping of orbits by approximate or exact local cost extrema, as a consequence of the highly complex cost landscape. Efficient algorithms have been devised to compensate for these obstacles in various scenarios, but from a mathematically rigorous perspective, there is at present insufficient insight into the conceptual and geometric core of the problem.
Our main goal here is to develop a rigorous mathematical understanding of some crucial dynamical systems aspects of supervised learning in DL networks; we do not focus on the development of efficient algorithms, or exploration of specific application cases. The definition of a gradient vector field necessitates the choice of a Riemannian metric structure, a general fact of differential geometry. In the standard approach to DL, the gradient of the cost function is determined using the Euclidean metric on the space of parameters (weights and biases); this choice is ad hoc, and not motivated by geometric reasons. As explained in (1.25), below, this defines a gradient flow in the output layer of the DL network with respect to a highly nontrivial induced metric determined by the Neural Tangent Kernel (NTK), [15]; the latter is the reason for the trapping of orbits.
We exploit the arbitrariness of choice in the Riemannian metric relative to which the gradient flow can be defined, both on the side of the parameter space, and in the output layer of the DL network. Choosing the gradient flow with respect to the Euclidean metric in the output layer of the DL network (instead of picking it on the space of parameters), we introduce two modified versions of the gradient descent flow in the parameter space, one adapted for the overparametrized setting, and the other for the underparametrized setting. Both have a natural invariant geometric meaning, in terms of the pullback vector bundle structure in the overparametrized, and the pushforward vector bundle structure in the underparametrized setting. In the output layer, the Euclidean structure greatly simplifies the cost landscape, and trivializes the gradient flow.
In the overparametrized case, we prove in Theorem 2.1 that, provided a rank condition holds, all orbits of the modified gradient flow drive the ${\mathcal{L}}^{2}$ cost to its global minimum at a uniform exponential convergence rate; this yields an a priori stopping time for any prescribed proximity to the global ${\mathcal{L}}^{2}$ cost minimum. In Section 2.4, we point out relations of the latter to sub-Riemannian geometry. It is known that invertibility of the NTK implies that zero loss is reachable, [17, 19]. However, trapping of orbits by approximate local minima of the standard gradient descent flow can occur; this is addressed in Section 3. In Section 4, we generalize the above framework to the situation in which the rank condition does not hold; in particular, we show that local equilibria can only exist if a rank loss occurs, and that generically, they are not isolated points, but elements of a critical submanifold of parameter space; related results for global minima of overparametrized networks were established in [9]. In the underparametrized situation, we show in Theorem 5.1 that the analogous version of the modified gradient flow is mapped to a geometrically natural constrained gradient flow. In Section 6, we show that in the borderline case separating the strictly overparametrized and underparametrized situations, the two modified gradient flows coincide. A key objective in this work is to elucidate and exploit the link between gradient flows in the parameter space and in the output layer.
In our previous work [6] with P. Muñoz Ewald, we provided an explicit construction of globally minimizing weights and biases for the ${\mathcal{L}}^{2}$ cost in underparametrized ReLU DL networks. In [7], we proved that those global minimizers can generically not be approximated via the gradient descent flow algorithm; this fact remains valid for the modified gradient descent flow obtained in this paper for the underparametrized case. Some key aspects of the analysis presented in the work at hand are related to the approach in [4] to classical Hamiltonian systems with non-holonomic constraints; see also [1, 3] for related background on sub-Riemannian geometry, and [10, 11] on dynamical systems and control theory aspects of machine learning. We refer to [2, 14, 12, 16, 18] and the references therein for some contextually related background on neural networks.
1.1. Definition of the mathematical model
We consider a DL network in the context of supervised learning, with training inputs given by $x_{j}^{(0)}\in{\mathbb{R}}^{M}$, $j=1,\dots,N$ where we refer to ${\mathbb{R}}^{M}$ as the input space. We assume that the outputs are given by $y_{\ell}\in{\mathbb{R}}^{Q}$, $\ell=1,\dots,Q$. We introduce the map
$\omega:\{1,\dots,N\}\rightarrow\{1,\dots,Q\}\,,$ (1.1)
which assigns the output label $\omega(j)$ to the $j$-th input label; that is, $x_{j}^{(0)}$ corresponds to the output $y_{\omega(j)}$. We write
${\underline{y}}_{\omega}:=(y_{\omega(1)}^{T},\dots,y_{\omega(N)}^{T})^{T}\in{\mathbb{R}}^{NQ}$ (1.2)
where $A^{T}$ is the transpose of the matrix $A$. We denote by $N_{i}$ the number of training inputs belonging to the output vector $y_{i}$, $i=1,\dots,Q$.
We assume that the DL network contains $L$ hidden layers; the $\ell$-th layer is defined on ${\mathbb{R}}^{M_{\ell}}$, and recursively determined by
$x_{j}^{(\ell)}=\sigma(W_{\ell}x_{j}^{(\ell-1)}+b_{\ell})\;\;\in{\mathbb{R}}^{M_{\ell}}$ (1.3)
via the weight matrix $W_{\ell}\in{\mathbb{R}}^{M_{\ell}\times M_{\ell-1}}$, bias vector $b_{\ell}\in{\mathbb{R}}^{M_{\ell}}$, and activation function $\sigma$. We assume that the output layer
$x_{j}^{(L+1)}=W_{L+1}x_{j}^{(L)}+b_{L+1}\;\;\in{\mathbb{R}}^{Q}$ (1.4)
contains no activation function, and that $\sigma$ has a Lipschitz continuous derivative (non-smooth activations like ReLU can be accommodated by a local mollification).
We let the vector of parameters $\underline{\theta}\in{\mathbb{R}}^{K}$ enlist all components of the weights $W_{\ell}$ and biases $b_{\ell}$, $\ell=1,\dots,L+1$, including those in the output layer, so that
$K=\sum_{\ell=1}^{L+1}(M_{\ell}M_{\ell-1}+M_{\ell})$ (1.5)
where $M_{0}\equiv M$ accounts for the input layer.
In the output layer, we define
$x_{j}[\underline{\theta}]:=x_{j}^{(L+1)}\in{\mathbb{R}}^{Q}$ (1.6)
for notational brevity, and obtain the ${\mathcal{L}}^{2}$ cost as
${\mathcal{C}}[{\underline{x}}[\underline{\theta}]]$ $=$ $\frac{1}{2N}\big{|}{\underline{x}}[\underline{\theta}]-{\underline{y}}_{\omega}\big{|}_{{\mathbb{R}}^{QN}}^{2}$ (1.7)
$=$ $\frac{1}{2N}\sum_{j=1}^{N}|x_{j}[\underline{\theta}]-y_{\omega(j)}|_{{\mathbb{R}}^{Q}}^{2}\,,$
where
${\underline{x}}:=(x_{1}^{T},\dots,x_{N}^{T})^{T}\in{\mathbb{R}}^{QN}\,,$ (1.8)
and where $|\bullet|_{{\mathbb{R}}^{n}}$ denotes the Euclidean norm on ${\mathbb{R}}^{n}$.
1.2. The standard gradient descent flow.
The goal of the gradient descent algorithm is to find parameters $\underline{\theta}\in{\mathbb{R}}^{K}$ that minimize the cost function. The standard approach is based on the gradient flow
$\partial_{s}\underline{\theta}(s)=-\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;\;,\;\underline{\theta}(0)=\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}\,.$ (1.9)
The vector field $\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\bullet]]:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{K}$ is Lipschitz continuous if the derivative of the activation function $\sigma$ is Lipschitz. It then follows that the existence and uniqueness theorem for ordinary differential equations applies to (1.9). From
$\partial_{s}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=-\big{|}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\big{|}_{{\mathbb{R}}^{K}}^{2}\leq 0\,,$ (1.10)
the cost ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ is monotone decreasing in $s\in{\mathbb{R}}_{+}$, and since ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\geq 0$ is bounded below, the limit ${\mathcal{C}}_{*}=\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ exists for any orbit $\{\underline{\theta}(s)|s\in{\mathbb{R}}_{+}\}$, and depends on the initial data, ${\mathcal{C}}_{0}={\mathcal{C}}[{\underline{x}}[\underline{\theta}(0)]]$. In applications, the initial data $\underline{\theta}_{0}\in{\mathbb{R}}^{K}$ is often chosen at random.
The convergence of ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ implies that $\lim_{s\rightarrow\infty}|\partial_{s}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]|=0$, and hence, $\lim_{s\rightarrow\infty}|\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]|_{{\mathbb{R}}^{K}}=0$ due to (1.10). Thus, the basic goal is to find ${\mathcal{C}}_{*}=\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]={\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]$ where $\underline{\theta}_{*}$ is a critical point of the gradient flow (1.9), satisfying $0=-\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]$.
As discussed in [7], there are major challenges to this strategy. First of all, the cost always converges to a stationary value, but not necessarily to the global minimum. Typically, there exist approximate or exact local minima (where the norm of the gradient is small or zero), and identifying ones that yield a sufficiently well-trained DL network is mostly based on ad hoc methods.
Moreover, as $s\rightarrow\infty$, neither does $\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]={\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]$ imply that $\underline{\theta}(s)$ converges to $\underline{\theta}_{*}$, nor to any other element of $\{\underline{\theta}_{**}\in{\mathbb{R}}^{K}\,|\,{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{**}]]={\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\}$, nor that $\underline{\theta}(s)$ converges at all, without further assumptions on ${\mathcal{C}}[{\underline{x}}[\bullet]]$. Therefore, while ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ always converges to a stationary value of the cost function under the gradient descent flow (1.9), $\underline{\theta}(s)$ cannot generally be assumed to converge to a minimizer $\underline{\theta}_{*}$. A simple illustrative example is given in [7]. Therefore, it is a key problem to be able to estimate $s_{0}>0$ such that $\underline{\theta}(s_{0})$ determines a sufficiently well trained DL network, up to a prescribed precision.
As we will discuss in detail in the work at hand, a main source for these challenges is the definition of the gradient in (1.9) with respect to the Euclidean metric on ${\mathbb{R}}^{K}$; the choice of this particular Riemannian structure is ad hoc, and not motivated by the geometry of the problem. Choosing the Euclidean structure on the output layer of the DL network instead, we will obtain modified gradient flows for $\underline{\theta}(s)$ both in the overparametrized and underparametrized cases that are better adapted to the inherent geometry of the problem. In the overparametrized case, we will show that the modified gradient flow always drives the ${\mathcal{L}}^{2}$ cost to its global minimum, at a uniform exponential rate, provided that a rank condition holds.
1.3. Euclidean gradient flow in the output layer.
A basic fact that we will exploit extensively in this work is the circumstance that the cost (1.7) depends on the parameters $\underline{\theta}\in{\mathbb{R}}^{K}$ only via its dependence on ${\underline{x}}[\underline{\theta}]\in{\mathbb{R}}^{QN}$, the vector of outputs from the last layer of the DL network.
In preparation of the main subject of our discussion, we first introduce the following simple comparison model (which may at this point be considered as a priori unrelated to (1.10)), defined by the gradient flow
$\partial_{s}{\underline{x}}(s)$ $=$ $-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]$ (1.11)
${\underline{x}}(0)$ $=$ ${\underline{x}}^{(0)}\in{\mathbb{R}}^{QN}$
with $s\in{\mathbb{R}}_{+}$. The gradient here is defined with the Euclidean metric on ${\mathbb{R}}^{QN}$, and as will be explained in Section 1.4, this system plays a major role in this work because the modified gradient flows that we will introduce on the space of parameters ${\mathbb{R}}^{K}$ are mapped to it (or a constrained version of it). Component-wise, it corresponds to
$\partial_{s}(x_{j}(s)-y_{\omega(j)})$ $=$ $-\frac{1}{N}(x_{j}(s)-y_{\omega(j)})$ (1.12)
for all $j=1,\dots,N$, and its explicit solution is given by
$x_{j}(s)-y_{\omega(j)}=e^{-\frac{s}{N}}(x_{j}(0)-y_{\omega(j)})$ (1.13)
with initial data $x_{j}(0)\in{\mathbb{R}}^{QN}$, so that using the expression (1.7), we find
${\mathcal{C}}[{\underline{x}}(s)]=e^{-\frac{2s}{N}}{\mathcal{C}}[{\underline{x}}(0)]\,.$ (1.14)
Notably, the exponential convergence rates in (1.13) and (1.14) are uniform with respect to the respective initial data. From (1.13), we obtain
${\underline{x}}_{*}:=\lim_{s\rightarrow\infty}{\underline{x}}(s)={\underline{y}}_{\omega}\,.$ (1.15)
In particular, this is the unique global minimizer of the ${\mathcal{L}}^{2}$ cost, since
$\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}(s)]={\mathcal{C}}[{\underline{x}}_{*}]=0$ (1.16)
by continuity of ${\mathcal{C}}$. Uniqueness follow from the fact that ${\mathcal{C}}$ is strictly convex in the variable ${\underline{x}}-{\underline{y}}_{\omega}$.
1.4. Geometrically adapted gradient descent
Perhaps surprisingly at first, the simplicity of (1.20) and explicitness of its solutions can be exploited for the DL network. Here, we present two geometrically adapted modifications of (1.9), one for the underparametrized situation $K<QN$, and one for the overparametrized situation $K\geq QN$. Both will be naturally linked to the comparison model (1.11).
The cost ${\mathcal{C}}[{\underline{x}}[\underline{\theta}]]$ depends on the weights and biases $\underline{\theta}\in{\mathbb{R}}^{K}$ only through its dependence on the vector ${\underline{x}}[\underline{\theta}]\in{\mathbb{R}}^{QN}$ in the output layer, therefore defining
$D[\underline{\theta}]$ ::=:= $\Big{[}\frac{\partial x_{j}[\underline{\theta}]}{\partial\theta_{\ell}}\Big{]}_{j=1,\dots,N,\;\ell=1,\dots,K}$ (1.20)
$=$ $\left[\begin{array}[]{ccc}\frac{\partial x_{1}[\underline{\theta}]}{\partial\theta_{1}}&\cdots&\frac{\partial x_{1}[\underline{\theta}]}{\partial\theta_{K}}\\ \cdots&\cdots&\cdots\\ \frac{\partial x_{N}[\underline{\theta}]}{\partial\theta_{1}}&\cdots&\frac{\partial x_{N}[\underline{\theta}]}{\partial\theta_{K}}\end{array}\right]\;\;\;\in{\mathbb{R}}^{QN\times K}\,,$
we observe that
$\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]=D^{T}[\underline{\theta}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]\,.$ (1.21)
Hence, (1.9) can be written in the form
$\partial_{s}\underline{\theta}(s)=-D^{T}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;\;,\;\underline{\theta}(0)=\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}\,,$ (1.22)
where, for the ${\mathcal{L}}^{2}$ cost, we obtain the simple expression
$\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=\frac{1}{N}({\underline{x}}[\underline{\theta}(s)]-{\underline{y}}_{\omega})\,.$ (1.23)
From here on, we will interchangeably write
${\underline{x}}(s)\equiv{\underline{x}}[\underline{\theta}(s)]$ (1.24)
for notational brevity.
The standard gradient flow (1.22) implies that
$\partial_{s}{\underline{x}}(s)$ $=$ $-D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (1.25)
${\underline{x}}(0)$ $=$ ${\underline{x}}[\underline{\theta}_{0}]\in{\mathbb{R}}^{QN}$
The right hand side of (1.25) is the gradient vector field defined with respect to the Riemannian structure on ${\mathbb{R}}^{QN}$ with metric tensor $(DD^{T})^{-1}$. This is the pushforward metric on ${\mathbb{R}}^{QN}$ induced by the Euclidean metric on ${\mathbb{R}}^{K}$ associated to (1.22), induced by the map ${\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]$. The matrix $DD^{T}$ corresponds to the Neural Tangent Kernel (NTK), [15].
We will discuss this fact in more detail in Section 3. This induced metric is the geometric origin of the complicated cost landscape; it is known that if $DD^{T}$ is invertible, zero loss can be reached, [17, 19]. However, orbits can get trapped in approximate local minima (locations where the norm of the gradient is small) which typically occur in large quantities, and whose distribution is very difficult to control.
1.4.1. The overparametrized case
Instead of using the gradient flow corresponding to the Euclidean metric in ${\mathbb{R}}^{K}$, we will now introduce a modified gradient flow for the overparametrized DL network $K\geq QN$.
We assume that $D$ has full rank $QN$ in the region $U\subseteq{\mathbb{R}}^{K}$, and let
${\rm Pen}[D]=D^{T}(DD^{T})^{-1}\;\;\in{\mathbb{R}}^{K\times QN}$ (1.26)
denote the Penrose inverse of $D$ on $U$, also of rank $QN$. It generalizes the notion of the inverse of a matrix by way of
$D{\rm Pen}[D]={\bf 1}_{QN\times QN}\,,$ (1.27)
and
$P:={\rm Pen}[D]D=P^{2}=P^{T}$ (1.28)
is the projector onto the range of $D^{T}$, orthogonal with respect to the Euclidean metric on ${\mathbb{R}}^{K}$.
Then, we define the modified gradient flow on $U$,
$\partial_{s}\underline{\theta}(s)$ $=$ $-{\rm Pen}[D[\underline{\theta}(s)]]({\rm Pen}[D[\underline{\theta}(s)]])^{T}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\,.$ (1.29)
We note that $P\,{\rm Pen}[D]={\rm Pen}[D]D{\rm Pen}[D]={\rm Pen}[D]$ due to (1.27). Therefore, the right hand side of (1.29) is mapped by $P[\underline{\theta}(s)]$ to itself, and lies in the range of $P[\underline{\theta}(s)]$. Hence, (1.29) has the structure of a constrained dynamical system in ${\mathbb{R}}^{K}$.
Explicitly, (1.29) can be written as
$\partial_{s}\underline{\theta}(s)$ $=$ $-D^{T}[\underline{\theta}(s)](D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)])^{-2}D[\underline{\theta}(s)]\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}\,,$ (1.30)
and as will be shown in Theorem 2.1, the vector ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]\in{\mathbb{R}}^{QN}$ corresponding to the output layer of the DL network can be verified to satisfy
$\partial_{s}{\underline{x}}(s)=D[\underline{\theta}(s)]\partial_{s}\underline{\theta}(s)=-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]\,,$ (1.31)
using (1.4.1) and (1.21). That is, it precisely corresponds to the Euclidean gradient flow of the comparison system (1.11), and therefore exhibits the same uniform exponential convergence rates as in (1.13) and (1.14).
Therefore, (1.29) always drives the cost to the global minimum, for any initial data $\underline{\theta}(0)$ for which ${\rm rank}(D[\underline{\theta}(s)])=QN$ is maximal along the associated orbit $\underline{\theta}(s)$. The latter condition ensures that $D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\in{\mathbb{R}}^{QN\times QN}$ is invertible, so that the right hand side of (1.4.1) is well-defined. In this situation,
$\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=0$ (1.32)
holds
The geometric structure will be described in detail in Section 2; the map ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$ defines the pullback vector bundle ${\mathcal{V}}\subset T{\mathbb{R}}^{K}$, endowed with the pullback metric $h$ induced by the Euclidean structure on ${\mathbb{R}}^{QN}$. Then, the operator ${\rm Pen}[D]({\rm Pen}[D])^{T}\nabla_{\underline{\theta}}$ on the right hand side of (1.29) is the coordinate representation of the gradient ${\rm grad}_{h}$ on ${\mathcal{V}}$. This ensures that the vector field (1.29) is horizontal with respect to ${\mathcal{V}}$, and in particular, that its pushforward under ${\underline{x}}$ is the gradient vector field $-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}]$ on $T{\mathbb{R}}^{QN}$ with the Euclidean metric. As we will explain in Section 2.4, ${\mathcal{V}}$ is in general non-holonomic, and the triple $({\mathbb{R}}^{K},{\mathcal{V}},h)$ defines a sub-Riemannian manifold.
In Section 4, we present a differential-algebraic system that generalizes (1.4.1) to allow for a loss of rank of $D$.
1.4.2. The underparametrized case
For the underparametrized DL network with $K<QN$, the map ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]$ is an embedding. We introduce the modified gradient flow defined by
$\partial_{s}\underline{\theta}(s)$ $=$ $-(D^{T}[\underline{\theta}(s)]D[\underline{\theta}(s)])^{-1}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}\,.$ (1.33)
In a similar manner as in (1.31), one obtains that this is equivalent to the constrained dynamical system
$\partial_{s}{\underline{x}}(s)=-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]$ (1.34)
for ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]$, where
${\mathcal{P}}=D(D^{T}D)^{-1}D^{T}={\mathcal{P}}^{2}={\mathcal{P}}^{T}\,$ (1.35)
is the projector associated with the tangent bundle of the embedding ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$. This holds for any initial data $\underline{\theta}(0)$ for which ${\rm rank}(D[\underline{\theta}(s)])=K$ is maximal along the associated orbit $\underline{\theta}(s)$.
2. Geometrically adapted gradient flow for overparametrized DL
We will now give the precise statement of the results described in Section 1.4.1, for the overparametrized DL network with $K\geq QN$. We have the following main theorem.
Theorem 2.1.
Given an overparametrized Deep Learning network with $K\geq QN$, assume for $D$ as in (1.20) that
${\rm rank}(D)=QN$ (2.1)
is maximal in the region $U\subset{\mathbb{R}}^{K}$. Let
${\rm Pen}[D]:=D^{T}(DD^{T})^{-1}\;\;\in{\mathbb{R}}^{K\times QN}$ (2.2)
denote the Penrose inverse of $D$ in $U$, with
${\rm Pen}[D]D=P\;\;,\;\;D{\rm Pen}[D]={\bf 1}_{QN\times QN}$ (2.3)
where $P=P^{2}=P^{T}\in{\mathbb{R}}^{K\times K}$ is the orthoprojector onto the range of $D^{T}\in{\mathbb{R}}^{K\times QN}$.
Assume that $\underline{\theta}(s)\in U$, $s\in{\mathbb{R}}_{+}$, is a solution of the modified gradient flow
$\partial_{s}\underline{\theta}(s)$ $=$ $-{\rm Pen}[D[\underline{\theta}(s)]]({\rm Pen}[D[\underline{\theta}(s)]])^{T}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\in U\,.$ (2.4)
Then, the vector associated to the output layer ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]\in{\mathbb{R}}^{QN}$ satisfies
$\partial_{s}{\underline{x}}(s)=-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]\;\;,\;\;{\underline{x}}(0)={\underline{x}}[\underline{\theta}_{0}]\in{\mathbb{R}}^{QN}\,.$ (2.5)
In particular, for any arbitrary initial data $\underline{\theta}(0)=\underline{\theta}_{0}$ allowing for (2.1) to hold along the corresponding orbit $\underline{\theta}(s)$, $s\in{\mathbb{R}}_{+}$, the ${\mathcal{L}}^{2}$ cost converges to its global minimum,
${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=e^{-\frac{2s}{N}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(0)]]\;\;\;(\rightarrow 0\;{\rm as}\;s\rightarrow\infty)\,,$ (2.6)
and
${\underline{x}}[\underline{\theta}(s)]-{\underline{y}}_{\omega}=e^{-\frac{s}{N}}({\underline{x}}[\underline{\theta}(0)]-{\underline{y}}_{\omega})\;\;\;(\rightarrow 0\;{\rm as}\;s\rightarrow\infty)\,,$ (2.7)
at the same uniform exponential convergence rates as in (1.13) and (1.14).
If $\underline{\theta}(s)$ itself converges as $s\rightarrow\infty$, then $\underline{\theta}_{*}:=\lim_{s\rightarrow\infty}\underline{\theta}(s)$ satisfies
$\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]=\frac{1}{N}({\underline{x}}[\underline{\theta}_{*}]-{\underline{y}}_{\omega})=0\,,$ (2.8)
and $\underline{\theta}_{*}$ is a global minimizer, ${\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]=0$.
Proof.
Given that ${\rm rank}(D[\underline{\theta}])=QN$ is maximal for $\underline{\theta}\in U\subset{\mathbb{R}}^{K}$, it follows that $DD^{T}$ is invertible on $U$, so that (2.2) and the right hand side of (2.1) are well-defined. We have
$\partial_{s}{\underline{x}}(s)$ $=$ $D[\underline{\theta}(s)]\,\partial_{s}\underline{\theta}$ (2.9)
$=$ $-D[\underline{\theta}(s)]\underbrace{D^{T}[\underline{\theta}(s)](D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)])^{-2}D[\underline{\theta}(s)]}_{={\rm Pen}[D[\underline{\theta}(s)]]({\rm Pen}[D[\underline{\theta}(s)]])^{T}}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-\underbrace{(D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)])^{-1}D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]}_{={\bf 1}_{QN\times QN}}\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]\,.$
where we used (1.21) to pass from the second to the third line.
Thus, it follows from the discussion in Section 1.3 that if (2.1) is satisfied along the orbit $\underline{\theta}(s)$ for all $s\in{\mathbb{R}}_{+}$, then
$\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=0\,.$ (2.10)
Convexity of ${\mathcal{C}}[{\underline{x}}]$ in ${\underline{x}}-{\underline{y}}_{\omega}$ implies that $\lim_{s\rightarrow\infty}{\underline{x}}[\underline{\theta}(s)]={\underline{y}}_{\omega}$ converges. The convergence rates are as in (1.13) and (1.14), because (2.9) is the same dynamical system as (1.11).
If $\underline{\theta}_{*}=\lim_{s\rightarrow\infty}\underline{\theta}(s)$ exists, then
$\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}[(s)]]]={\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]=0\,,$ (2.11)
by continuity of ${\mathcal{C}}$ and of ${\underline{x}}[\bullet]$. Then, (2.8) follows from ${\mathcal{C}}[{\underline{x}}]=\frac{N}{2}|\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}]|^{2}$. ∎
2.1. Rank condition on $DD^{T}$
The rank of $DD^{T}$ plays a major role in the above considerations. In [15], a detailed discussion is provided about the fact that in the infinite width limit of a deep network, the NTK $DD^{T}$ is positive definite, and thus of full rank. Without the assumption of the infinite width limit, it is a subtle and challenging problem to determine sufficient conditions for full rank. In [8], the NTK is proven to have full rank on an open subset of parameter space ${\mathbb{R}}^{K}$ in overparametrized deep networks with sufficient width and for generic training data; but the question whether this property is preserved along the entire gradient descent trajectory remains to be investigated in future work.
2.2. A priori stopping condition
While for the geometrically adapted gradient flow (2.1), the ${\mathcal{L}}^{2}$ cost always converges at an exponential rate to its global minimum, as long as the rank condition (2.1) holds, orbits $\underline{\theta}(s)$, $s\in{\mathbb{R}}_{+}$, do not necessarily converge. However, because of the exponential convergence of the cost along any orbit $\underline{\theta}(s)$ as given in (1.14), we obtain the following sharp stopping time.
Theorem 2.2.
Assume that ${\rm rank}(DD^{T})=QN$ is maximal along the orbit $\underline{\theta}(s)$ of (2.1). Given $\epsilon>0$, let
$s_{\epsilon,\underline{\theta}(0)}:=\inf_{s\geq 0}\big{\{}s\in{\mathbb{R}}_{+}\,\big{|}\,{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\leq\epsilon\,\big{\}}$ (2.12)
denote the earliest time at which the cost is $\leq\epsilon$. Then,
$s_{\epsilon,\underline{\theta}(0)}=\frac{N}{2}\log\frac{{\mathcal{C}}[{\underline{x}}[\underline{\theta}(0)]]}{\epsilon}\,.$ (2.13)
Proof.
We have
$s_{\epsilon,\underline{\theta}(0)}=\inf_{s\geq 0}\big{\{}s\in{\mathbb{R}}_{+}\,\big{|}\,{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=e^{-\frac{2s}{N}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(0)]]\leq\epsilon\,\big{\}}\,,$ (2.14)
due to (2.6). This immediately implies that
$s_{\epsilon,\underline{\theta}(0)}=\frac{N}{2}\log\frac{{\mathcal{C}}[{\underline{x}}[\underline{\theta}(0)]]}{\epsilon}\,,$ (2.15)
as claimed. ∎
That is, for any given $\epsilon>0$, we obtain a precise stopping time for the modified gradient flow at
$s_{stop}\geq s_{\epsilon,\underline{\theta}(0)}\,.$ (2.16)
The DL network trained with weights and biases $\underline{\theta}(s_{stop})$ then produces an ${\mathcal{L}}^{2}$ cost ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s_{stop})]]\leq\epsilon$. The key advantage of this circumstance is that the minimum stopping time $s_{\epsilon,\underline{\theta}(0)}$ can be determined a priori; except for the initial value of the ${\mathcal{L}}^{2}$ cost, it is independent of any orbit satisfying (2.1). Therefore, a precise estimate on the training time based on the modified gradient descent flow can be given at the onset, provided that the rank condition ${\rm rank}(D)=QN$ in (2.1) is satisfied along the orbit.
2.3. Gradient flow with respect to pullback bundle metric.
We first describe the geometric setting for the general case, before specifying it for the situation in discussion. Let ${\mathcal{M}}$ and ${\mathcal{N}}$ be manifolds of dimensions $k={\rm dim}({\mathcal{M}})>n={\rm dim}({\mathcal{N}})$, and assume that ${\mathcal{N}}$ is endowed with a Riemannian structure $({\mathcal{N}},g)$ with metric $g$. Let $T{\mathcal{M}}$ and $T{\mathcal{N}}$ denote the respective tangent bundles, and assume that $f:{\mathcal{M}}\rightarrow{\mathcal{N}}$ is a smooth surjection. Within this section, ”smooth” will mean sufficiently regular, which will mostly be $C^{1}$. We denote the corresponding pullback map by $f^{*}$, and the pushforward map by $f_{*}$.
Let $\Gamma({\mathcal{M}})$ denote the set of smooth sections (vector fields) $V:{\mathcal{M}}\rightarrow T{\mathcal{M}}$, and similarly for $\Gamma({\mathcal{N}})$. We denote by ${\mathcal{V}}\subset T{\mathcal{M}}$ the pullback vector bundle $f^{*}T{\mathcal{N}}$, and the space of smooth sections $V:{\mathcal{M}}\rightarrow{\mathcal{V}}$ by $\Gamma({\mathcal{V}})$. That is, for every $z\in{\mathcal{M}}$, the fiber ${\mathcal{V}}_{z}\subset T_{z}{\mathcal{M}}$ is spanned by $f^{*}w$ for vectors $w\in T_{f(z)}{\mathcal{N}}$. In particular, we may define the pullback metric $h$ on ${\mathcal{V}}$ by way of
$h(V,W)=g(f_{*}V,f_{*}W)$ (2.17)
for sections $V,W\in\Gamma({\mathcal{V}})$. Then, we can define the gradient ${\rm grad}_{h}$ associated to $({\mathcal{V}},h)$ by way of
$d{\mathcal{F}}(V)=h(V,{\rm grad}_{h}{\mathcal{F}})$ (2.18)
for all $V\in\Gamma({\mathcal{V}})$, and any smooth ${\mathcal{F}}:{\mathcal{M}}\rightarrow{\mathbb{R}}$, where $d{\mathcal{F}}$ denotes its exterior derivative.
In our concrete situation, ${\mathcal{M}}={\mathbb{R}}^{K}$, ${\mathcal{N}}\subseteq{\mathbb{R}}^{QN}$ open, so that $k=K$ and $n=QN$. Moreover, the coordinate representation of the map $f$ is ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$ and that of the pushforward $f_{*}$ applied to vector fields is the Jacobian matrix $[Df^{\alpha}_{\beta}]=D[\underline{\theta}]$ defined in (1.20). For the pullback vector bundle ${\mathcal{V}}$ under ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, we have that the fiber ${\mathcal{V}}_{\underline{\theta}}$ is given by the range of $D^{T}[\underline{\theta}]$. We claim that the modified gradient flow (2.19) is the gradient flow with respect to the bundle metric $h$ associated to ${\mathcal{V}}$.
Proposition 2.3.
Let $h$ denote the pullback metric on the vector bundle ${\mathcal{V}}$ induced by the Euclidean metric on $T{\mathbb{R}}^{QN}$, as in (2.20). Let ${\rm grad}_{h}$ denote the gradient with respect to $h$ as in (2.25). Then, writing $\widetilde{\mathcal{C}}[\underline{\theta}]:={\mathcal{C}}[{\underline{x}}[\underline{\theta}]]$ for the ${\mathcal{L}}^{2}$ cost function, the geometrically invariant expression for (2.1) is given by
$\partial_{s}\underline{\theta}(s)=-{\rm grad}_{h}\widetilde{\mathcal{C}}[\underline{\theta}(s)]\,.$ (2.19)
In particular, the gradient vector field in (2.19) is a section of ${\mathcal{V}}$, and thus horizontal with respect to ${\mathcal{V}}$.
Proof.
In local coordinates, we let $g_{\alpha,\alpha^{\prime}}$ with $\alpha,\alpha^{\prime}=1,\dots,n$ denote the matrix components of $g$, and $\sum_{\beta}Df^{\alpha}_{\beta}V^{\beta}$ with $\alpha=1,\dots,k$, $\beta=1,\dots,K$, the coordinate representation of $f_{*}V$ where $V=\sum_{\beta}V^{\beta}\partial_{\beta}$, and $Df^{\alpha}_{\beta}$ are the components of the Jacobian matrix of $f$. Then,
$h(V,W)=\sum_{\alpha,\alpha^{\prime},\beta,\beta^{\prime}}g_{\alpha,\alpha^{\prime}}Df^{\alpha}_{\beta}Df^{\alpha^{\prime}}_{\beta^{\prime}}V^{\beta}W^{\beta^{\prime}}\,,$ (2.20)
so that (2.18) is equivalent to
$\sum_{\beta}V^{\beta}\partial_{\beta}{\mathcal{F}}=\sum_{\alpha,\alpha^{\prime},\beta,\beta^{\prime}}g_{\alpha,\alpha^{\prime}}Df^{\alpha}_{\beta}Df^{\alpha^{\prime}}_{\beta^{\prime}}V^{\beta}({\rm grad}_{h}{\mathcal{F}})^{\beta^{\prime}}$ (2.21)
for all $V\in\Gamma({\mathcal{V}})$, from $d{\mathcal{F}}(V)=\sum_{\beta}V^{\beta}\partial_{\beta}{\mathcal{F}}$.
The Riemannian structure on ${\mathcal{N}}\subseteq{\mathbb{R}}^{QN}$ is given by the Euclidean metric, $g_{\alpha,\alpha^{\prime}}=\delta_{\alpha,\alpha^{\prime}}$ (the Kronecker delta) in Euclidean coordinates, so that (5.16) has the form
${\underline{V}}^{T}\nabla_{\underline{\theta}}{\mathcal{F}}={\underline{V}}^{T}D^{T}[\underline{\theta}]D[\underline{\theta}]\;{\rm grad}_{h}{\mathcal{F}}\,,$ (2.22)
for all ${\underline{V}}=(V^{1},\dots,V^{K})^{T}\in{\rm range}(D^{T}[\underline{\theta}])$. Thus, this is equivalent to
$P[\underline{\theta}]\nabla_{\underline{\theta}}{\mathcal{F}}=P[\underline{\theta}]D^{T}[\underline{\theta}]D[\underline{\theta}]\;{\rm grad}_{h}{\mathcal{F}}$ (2.23)
where $P=P^{2}=P^{T}\in{\mathbb{R}}^{K\times K}$ is the orthoprojector onto the range of $D^{T}$, as defined in (2.3). As noted above, the range of $D^{T}[\underline{\theta}]$ is, by construction, equivalent to the fiber ${\mathcal{V}}_{\underline{\theta}}$. Applying the Penrose inverse of $D^{T}[\underline{\theta}]$ from the left, we obtain
$({\rm Pen}[D[\underline{\theta}]])^{T}P[\underline{\theta}]\nabla_{\underline{\theta}}{\mathcal{F}}=({\rm Pen}[D[\underline{\theta}]])^{T}\nabla_{\underline{\theta}}{\mathcal{F}}=D[\underline{\theta}]\;{\rm grad}_{h}{\mathcal{F}}\,,$ (2.24)
and subsequently applying the Penrose inverse of $D[\underline{\theta}]$ from the left,
${\rm grad}_{h}{\mathcal{F}}={\rm Pen}[D[\underline{\theta}]]({\rm Pen}[D[\underline{\theta}]])^{T}\nabla_{\underline{\theta}}{\mathcal{F}}\,.$ (2.25)
This is the expression for the gradient associated to the pullback metric $h$ on the vector bundle ${\mathcal{V}}$, in local coordinates. Notably, $P{\rm grad}_{h}{\mathcal{F}}={\rm grad}_{h}{\mathcal{F}}$, and $P^{\perp}{\rm grad}_{h}{\mathcal{F}}=0$; that is, ${\rm grad}_{h}{\mathcal{F}}\in\Gamma({\mathcal{V}})$ as defined here is a section of ${\mathcal{V}}$. This establishes the claim. ∎
2.4. Relationship to sub-Riemannian geometry
It is natural to ask whether the vector bundle ${\mathcal{V}}$ is integrable (holonomic), i.e., whether ${\mathbb{R}}^{K}$ foliates into submanifolds whose tangent spaces coincide with the fibers of ${\mathcal{V}}$. This holds if and only if the Frobenius condition is satisfied,
${\mathcal{V}}\;{\rm integrable}\;\Leftrightarrow\;[V,W]\in\Gamma({\mathcal{V}})\;\;,\;\;\forall V,W\in\Gamma({\mathcal{V}})\,,$ (2.26)
that is, the Lie bracket of any pair of sections of ${\mathcal{V}}$ is a section of ${\mathcal{V}}$. To our knowledge, this is not generally the case for the situation under consideration.
To address the latter point, we observe that ${\mathcal{V}}$ is the normal bundle to the foliation
${\mathbb{R}}^{K}=\bigcup_{x\in{\mathbb{R}}^{QN}}f^{-1}(x)\;\;{\rm where}\;\;f:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}\;,\;\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]\,.$ (2.27)
That is, for any $\underline{\theta}\in f^{-1}(x)\subset{\mathbb{R}}^{K}$, ${\mathcal{V}}_{\underline{\theta}}$ is normal to $T_{\underline{\theta}}f^{-1}(x)$. However, this does not imply integrability of ${\mathcal{V}}$. As an example of a foliation whose normal bundle is non-integrable, we may consider any smooth, non-integrable distribution $\widetilde{\mathcal{V}}\subset T{\mathbb{R}}^{K}$ of fiber rank $K-1$. Then, $\widetilde{{\mathcal{V}}}$ is normal to a vector bundle of fiber rank 1. The latter is always integrable (the Frobenius condition is trivially satisfied, as the Lie bracket of any section with itself is zero), and thus generates a foliation of ${\mathbb{R}}^{K}$ into 1-dimensional curves. Accordingly, $\widetilde{{\mathcal{V}}}$ is a non-integrable normal bundle to a foliation.
In case of ${\mathcal{V}}$ being non-integrable, (2.19) is obtained by imposing non-holonomic constraints on the standard gradient flow (1.9); the triple
$({\mathbb{R}}^{K},{\mathcal{V}},h)$ (2.28)
then defines a sub-Riemannian manifold. For some context on non-holonomic systems and sub-Riemannian geometry, we refer to [1, 3].
3. Comparison with the standard overparametrized gradient flow
In comparison, we consider the standard gradient flow (1.22) with the Euclidean metric in the parameter space ${\mathbb{R}}^{K}$,
$\partial_{s}{\underline{x}}(s)$ $=$ $-D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]$ (3.1)
${\underline{x}}(0)$ $=$ ${\underline{x}}[\underline{\theta}_{0}]\;\in{\mathbb{R}}^{QN}\,,$
in the overparametrized scenario, $K>QN$, where ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]$. The operator $DD^{T}\nabla_{{\underline{x}}}$ is the gradient associated to the Riemannian structure with metric tensor $(DD^{T})^{-1}|_{f^{-1}({\underline{x}})}$ (assuming invertibility) on ${\mathbb{R}}^{QN}$ where $f:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]$. This is the pushforward metric on ${\mathbb{R}}^{QN}$ induced by the Euclidean metric on ${\mathbb{R}}^{K}$ associated to (1.22), under the map $f$.
3.1. Trapping of orbits
This pushforward metric is the source of the complicated cost landscape encountered for the standard gradient flow, and the emergence of a multitude of approximate local minima (where the norm of $DD^{T}\nabla_{{\underline{x}}}{\mathcal{C}}$ in (3.1) is small) which can trap its orbits.
We write the standard gradient descent flow (1.9) in the form
$\partial_{s}\underline{\theta}(s)$ $=$ $-D^{T}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (3.2)
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}$
using (1.21).
It is known that invertibility of $DD^{T}$ implies that zero loss can be reached, [17, 19]. However, the main issue here arises from the fact that a priori, there exists no uniform lower bound on the NTK $DD^{T}$, that is,
$\nexists\lambda>0\,:\;\;(DD^{T})[\underline{\theta}]>\lambda\;,\;\forall\underline{\theta}\in{\mathbb{R}}^{K}\,.$ (3.3)
Therefore, there is no mechanism precluding
$\inf_{\underline{\theta}\in{\mathbb{R}}^{K}}\inf_{v\in{\mathbb{R}}^{QN},|v|=1}|\,(DD^{T})[\underline{\theta}]\,v\,|$ (3.4)
from being arbitrarily small, even if $DD^{T}>0$. As a consequence, there may exist regions that ”trap” orbits of (3.2) for a long time in the following sense, even if ${\rm rank}(D^{T})=QN$ is maximal, so that $DD^{T}>0$ holds.
Proposition 3.1.
Assume that there exists a region $U\subset{\mathbb{R}}^{K}$ such that
$0<|D^{T}[\underline{\theta}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]|<\epsilon|\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]|\ll 1\;\;,\;\forall\underline{\theta}\in U\,,$ (3.5)
for some $0<\epsilon\ll 1$. Moreover, assume that the orbit of the gradient flow (3.2) intersects $U$, that is, $\underline{\theta}(s)\in U$ for $s\in I\subseteq{\mathbb{R}}_{+}$, and let
$L_{U}:=\big{|}\,\{\underline{\theta}(s)|s\in I\}\cap U\,\big{|}$ (3.6)
denote the arc length of its intersection with $U$. Then, with $s_{0}:=\inf I$,
$|I|\,\geq\,\frac{N\,L_{U}}{|{\underline{x}}[\underline{\theta}(s_{0})]-{\underline{y}}_{\omega}|}\,\frac{1}{\epsilon}$ (3.7)
is a lower bound on the time during which the orbit is trapped inside $U$.
Proof.
Clearly,
$\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}]=\frac{1}{N}({\underline{x}}-{\underline{y}}_{\omega})\;\in{\mathbb{R}}^{QN}$ (3.8)
is zero if and only if ${\underline{x}}={\underline{y}}_{\omega}$, which globally minimizes the cost, due to (1.7).
Then, we consider the arc length of the intersection of the orbit with $U$,
$L_{U}$ $=$ |$\big{|}\,\{\underline{\theta}(s)|s\in{\mathbb{R}}_{+}\}\cap U\,\big{|}$ (3.9)
[…]
$\leq$ $|I|\epsilon\;\sup_{s\in I}|\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]|$
$=$ $|I|\epsilon\;\Big{(}\frac{2}{N}\sup_{s\in I}|{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]|\Big{)}^{\frac{1}{2}}$
$=$ $|I|\epsilon\;\Big{(}\frac{2}{N}|{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s_{0})]]|\Big{)}^{\frac{1}{2}}$
$=$ $\frac{|I|\epsilon}{N}\;|{\underline{x}}[\underline{\theta}(s_{0})]-{\underline{y}}_{\omega}|$
where $s_{0}=\inf I$ is the initial time of entry of the orbit into $U$. Here, we used monotone decrease of the cost along the orbit to pass to the fifth line, and recalled (1.7) to pass to the last line. The claim follows. ∎
In the context of computationally determining orbits of the standard gradient descent flow, ”trapping regions” of this type might turn out to be difficult to distinguish from true (degenerate) local extrema where $\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]=0$. Within the bounds of numerical precision for a given implementation, they might indeed be indistinguishable.
3.2. Local extrema and rank loss
In contrast to the situation addressed in Proposition 3.1 which assumes full rank of $D$, we note that true local extrema of (3.2) outside of the global cost minimum necessitate a rank loss, ${\rm rank}(D)<QN$, as shown in the following lemma.
Lemma 3.2.
Any equilibrium solution $\underline{\theta}_{*}$ of the standard gradient flow (3.2) that is not a global minimizer of the cost, i.e., $\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]=0$ but $\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\neq 0$, must satisfy
${\rm rank}(D[\underline{\theta}_{*}])<QN\;\;{\rm and}\;\;\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\in\ker(D^{T}[\underline{\theta}_{*}])\,.$ (3.10)
Proof.
Let us first assume that ${\rm rank}(D[\underline{\theta}(s)])=QN$ along an orbit of (3.2). If this orbit contains an isolated local equilibrium point $\underline{\theta}_{*}$ of (3.2), then it must stem from
$0\neq\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\in\ker(D^{T}[\underline{\theta}_{*}])\,.$ (3.11)
However, because the domain of $D^{T}[\underline{\theta}(s)])$ is ${\mathbb{R}}^{QN}$, the assumption ${\rm rank}(D[\underline{\theta}(s)])=QN$ implies that its kernel is trivial. Therefore, (3.11) is not possible.
Therefore, existence of an isolated local equilibrium point $\underline{\theta}_{*}$ of (3.2), which is not a global minimimum, requires that both ${\rm rank}(D[\underline{\theta}_{*}])<QN$ and (3.11) hold.
4. Rank loss in overparametrized Deep Learning networks
In this section, we further investigate the standard and geometrically adapted DL networks in the overparametrized case, $K>QN$, but in the situation where a rank loss ${\rm rank}(D)<QN$ is allowed to occur.
4.1. Rank loss and constrained dynamics
If a rank loss occurs, ${\rm rank}(D^{T})<QN$ in a set $U^{\prime}\subset{\mathbb{R}}^{K}$, it follows that on $U^{\prime}$, the system is effectively underparametrized. To elucidate this point, we diagonalize the positive semi-definite matrix
$DD^{T}=R^{T}\Lambda R\;\;,\;\underline{\theta}\in U^{\prime}$ (4.1)
in $U^{\prime}$, where $R:U^{\prime}\rightarrow SO(QN)$, and $\Lambda$ maps $U^{\prime}$ to the set of positive semi-definite diagonal matrices. Let $\widetilde{\Lambda^{-1}}$ denote its generalized inverse, so that ${\mathcal{P}}_{\Lambda}:=\Lambda\widetilde{\Lambda^{-1}}=\widetilde{\Lambda^{-1}}\Lambda$ is the diagonal projector obtained from replacing all non-zero entries of $\Lambda$ by 1, while all other entries are zero. Accordingly,
${\mathcal{P}}:=R^{T}{\mathcal{P}}_{\Lambda}R\,=\,{\mathcal{P}}^{2}\,=\,{\mathcal{P}}^{T}$ (4.2)
is the orthoprojector (relative to the Euclidean metric on ${\mathbb{R}}^{QN}$) onto the range of $DD^{T}$ on $U^{\prime}$. In particular, $DD^{T}={\mathcal{P}}DD^{T}{\mathcal{P}}$ has full rank on the range of ${\mathcal{P}}$, and
$({\mathcal{P}}DD^{T}{\mathcal{P}})^{-1}|_{{\rm range}({\mathcal{P}})}=R^{T}\widetilde{\Lambda^{-1}}R\,.$ (4.3)
We also let ${\mathcal{P}}^{\perp}:={\bf 1}_{QN\times QN}-{\mathcal{P}}$.
4.1.1. Rank loss in the standard gradient flow
Given the above circumstances, the standard gradient descent flow yields
$\partial_{s}{\underline{x}}(s)$ $=$ $-{\mathcal{P}}[\underline{\theta}(s)]D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]\,,$ (4.4)
for ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]\in{\mathbb{R}}^{QN}$, $\underline{\theta}(s)\in U^{\prime}$.
Thus, it has the structure of a constrained dynamical system in ${\mathbb{R}}^{QN}$ relative to (3.1), due to the projection of the vector field to the range of ${\mathcal{P}}$. This is analogous to the situation encountered in underparametrized DL networks addressed in [7], and in the next section.
4.1.2. Rank loss in the geometrically adapted gradient flow
The definition of the geometrically adapted gradient flow given in Theorem 2.1 needs to be modified in case a rank loss, ${\rm rank}(D)<QN$, is admitted. In fact, we instead define the following differential-algebraic system; for background on the efficient computational solution of differential-algebraic systems, see for instance [13].
Theorem 4.1.
Assume that ${\rm rank}(D)\leq QN$, and that the orbit $\underline{\theta}(s)$ solves the differential-algebraic system
$\partial_{s}\underline{\theta}(s)$ $=$ $D^{T}[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]$ (4.5)
$\Psi[\underline{\theta}(s)]$ $=$ ${\rm argmin}_{\Psi}\{\;|D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\Psi+\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;|_{{\mathbb{R}}^{QN}}^{2}\;\}$
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\in{\mathbb{R}}^{K}\,.$
That is, $\Psi[\underline{\theta}(s)]$ solves
$D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\Psi=-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;\;+\;{\rm minimal\;error\;in}\;L^{2}\,$ (4.6)
via least square optimization.
Then, if $rank(D)=QN$ is maximal, (4.5) is equivalent to the geometrically adapted gradient flow in Theorem 2.1.
If ${\rm rank}(D)<QN$, then ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]$ solves
$\partial_{s}{\underline{x}}(s)$ $=$ $-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (4.7)
${\underline{x}}(0)$ $=$ ${\underline{x}}[\underline{\theta}_{0}]\in{\mathbb{R}}^{QN}\,,$
where the projector ${\mathcal{P}}$ onto the range of $DD^{T}$ is defined in (4.2).
Proof.
Evidently, if ${\rm rank}(D)=QN$ is maximal, then $DD^{T}$ is invertible, and (4.5) reduces to the definition in Theorem 2.1. On the other hand, if ${\rm rank}(D)<QN$, then it follows from
$|D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\Psi+\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;|_{{\mathbb{R}}^{QN}}^{2}$
$=$ $|{\mathcal{P}}[\underline{\theta}(s)]D[\underline{\theta}(s)]D^{T}[\underline{\theta}(s)]\Psi+{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;|_{{\mathbb{R}}^{QN}}^{2}$
$\hskip 113.81102pt+|{\mathcal{P}}^{\perp}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\;|_{{\mathbb{R}}^{QN}}^{2}$
that the right hand side is minimized when the first term on the rhs is zero, whence
$({\mathcal{P}}DD^{T}{\mathcal{P}})[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]=-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (4.9)
where we used that ${\mathcal{P}}DD^{T}=DD^{T}=DD^{T}{\mathcal{P}}$. Because ${\mathcal{P}}DD^{T}{\mathcal{P}}$ has full rank on the range of ${\mathcal{P}}$, we find that
$\Psi[\underline{\theta}(s)]=-({\mathcal{P}}DD^{T}{\mathcal{P}})^{-1}|_{{\rm range}({\mathcal{P}})}[\underline{\theta}(s)]{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (4.10)
We thus find for ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]$ that
$\partial_{s}{\underline{x}}(s)$ $=$ $D[\underline{\theta}(s)]\partial_{s}\underline{\theta}(s)$ (4.11)
$=$ $(DD^{T})[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]$
$=$ $-\big{(}\,({\mathcal{P}}DD^{T}{\mathcal{P}})({\mathcal{P}}DD^{T}{\mathcal{P}})^{-1}|_{{\rm range}({\mathcal{P}})}\,\big{)}[\underline{\theta}(s)]{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\,.$
This proves the claim. ∎
Theorem 4.2.
A point $\underline{\theta}_{*}\in{\mathbb{R}}^{K}$ is an equilibrium of the standard gradient flow (3.2) if and only if it is an equilibrium either of the geometrically adapted gradient flow (4.5) if $D$ does not have full rank, or of (2.1) if $D$ has full rank.
Assume that the activation function $\sigma$ is smooth. If ${\rm rank}({\mathcal{P}})=r<QN$ in an open neighborhood $U\subset{\mathbb{R}}^{K}$, then local equilibria of (3.2) and (4.5) which are not global are generically contained in an $(K-r)$-dimensional critical submanifold ${\mathcal{M}}_{crit}\subset U$, in the sense of Sard.
Proof.
If a rank loss occurs, ${\rm rank}(D^{T})<QN$, in a neighborhood $U^{\prime}\subset{\mathbb{R}}^{K}$, then an equilibrium of (4.4) and an equilibrium of (4.7) both require the same condition
$0=-{\mathcal{P}}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\;\;,\;\underline{\theta}_{*}\in U^{\prime}\,.$ (4.12)
For (4.4), this is because of $DD^{T}={\mathcal{P}}DD^{T}{\mathcal{P}}$ has full rank on the range of ${\mathcal{P}}$. Therefore, any equilibrium of (2.1) that is not a global minimizer satisfies (3.10).
Next, assume that ${\rm rank}({\mathcal{P}})=r<QN$ on an open neighborhood $U\subset{\mathbb{R}}^{K}$, and let $\underline{\theta}\in U$. Given that the activation function $\sigma$ is smooth, both $D[\underline{\theta}]$ and ${\underline{x}}[\underline{\theta}]$ depend smoothly on $\underline{\theta}$. Assume that the linearly independent family of smooth vector fields $V_{\alpha}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\alpha=1,\dots,r$, spans the range of ${\mathcal{P}}$; for instance, one could choose a suitable $r$-tuple of column vectors of $D$. Then, we define the smooth family of functions $g_{\alpha}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}$ obtained from the inner product
$g_{\alpha}[\underline{\theta}]$ ::=:= $\big{(}\,V_{\alpha}[\underline{\theta}]\,,\,{\mathcal{P}}[\underline{\theta}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]\,\big{)}_{{\mathbb{R}}^{QN}}$ (4.13)
$=$ $\big{(}\,V_{\alpha}[\underline{\theta}]\,,\,\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}]]\,\big{)}_{{\mathbb{R}}^{QN}}\;,\;\alpha=1,\dots,r\,.$
The set of equilibrium solutions in $U\subset{\mathbb{R}}^{K}$ is given by
${\mathcal{M}}_{crit}=U\cap\bigcap_{\alpha=1}^{r}g_{\alpha}^{-1}(0)\,.$ (4.14)
Then, by Sard’s theorem, ${\mathcal{M}}_{crit}\subset U\subset{\mathbb{R}}^{K}$ is generically a $(K-r)$-dimensional submanifold. ∎
We conclude that the standard gradient flow (3.2) and the geometrically adapted gradient flow (2.1) have the same set of equilibria. If those equilibria are not global, then they are generically not isolated points, but elements of a critical manifold. On the other hand, (2.1) exhibits a uniform exponential convergence rate whenever $DD^{T}$ has full rank, as opposed to (3.2) which provides no such a priori control. We also note that in [9], it is shown that global minima in overparametrized networks define critical submanifolds.
4.2. Perturbation away from local equilibrium
If an orbit converges towards (is trapped by) a local equilibrium $\underline{\theta}_{*}\in{\mathcal{M}}_{crit}$, one may wish to select an updated initial condition $\underline{\theta}_{0}^{\prime}$ to continue the gradient descent either based on the standard or geometrically adapted flow, that is, (4.4) or (4.7), respectively. A suitable choice is obtained from
$\underline{\theta}_{0}^{\prime}=\underline{\theta}_{*}+V_{\underline{\theta}_{*}}$ (4.15)
where ${\underline{x}}_{*}\in{\mathcal{M}}_{crit}$, and $V_{\bullet}:{\mathcal{M}}_{crit}\rightarrow N{\mathcal{M}}_{crit}$ is a section of the normal bundle of ${\mathcal{M}}_{crit}$. Since $N{\mathcal{M}}_{crit}$ is spanned by $\{\nabla_{\underline{\theta}}g_{\alpha}\big{|}_{{\mathcal{M}}_{crit}}\}_{\alpha=1}^{r}$, one can for example pick
$V_{\underline{\theta}_{*}}=\kappa_{\alpha}\nabla_{\underline{\theta}}g_{\alpha}\big{|}_{\underline{\theta}=\underline{\theta}_{*}}$ (4.16)
with $g_{\alpha}$ defined as in (4.13), and a suitable choice of $\alpha\in\{1,\dots,r\}$. The sign of the parameter $\kappa_{\alpha}\in{\mathbb{R}}$ is chosen such that the cost is decreasing,
${\mathcal{C}}[{\underline{x}}[\underline{\theta}_{0}^{\prime}]]<{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\,.$ (4.17)
More generally, one may choose a linear superposition of vectors of the form (4.16), for each of which one obtains (4.17).
5. Geometrically adapted gradient flow for underparametrized DL
In the underparametrized situation $K<QN$, we obtain the following theorem.
Theorem 5.1.
Assume that $K<QN$, and that
${\rm rank}(D[\underline{\theta}])=K$ (5.1)
is maximal in the region $\underline{\theta}\in U\subset{\mathbb{R}}^{K}$. Assume that $\underline{\theta}(s)\in U$, $s\in[0,T)\subseteq{\mathbb{R}}_{+}$, is an orbit of the modified gradient descent flow
$\partial_{s}\underline{\theta}(s)$ $=$ $-(D^{T}[\underline{\theta}(s)]D[\underline{\theta}(s)])^{-1}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (5.2)
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\;\in U\,.$
Then, ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]]$ is a solution to the constrained gradient flow defined by
$\partial_{s}{\underline{x}}(s)$ $=$ $-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ (5.3)
${\underline{x}}(0)$ $=$ ${\underline{x}}[\underline{\theta}_{0}]\in{\mathbb{R}}^{QN}$
where
${\mathcal{P}}=D(D^{T}D)^{-1}D^{T}\;\;\in{\mathbb{R}}^{QN\times QN}$ (5.4)
is the projector onto the range of $D$, orthogonal with respect to the Euclidean metric on ${\mathbb{R}}^{QN}$, satisfying ${\mathcal{P}}={\mathcal{P}}^{2}={\mathcal{P}}^{T}$. In particular,
${\mathcal{P}}^{\perp}[\underline{\theta}(s)]\partial_{s}{\underline{x}}(s)=0$ (5.5)
where ${\mathcal{P}}^{\perp}={\bf 1}_{QN\times QN}-{\mathcal{P}}$.
Moreover, any critical point $\underline{\theta}_{*}\in U$ satisfies
$0={\mathcal{P}}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\,,$ (5.6)
and the local extremum of the cost function at $\underline{\theta}_{*}$ is attained at
${\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]=\frac{N}{2}\big{|}{\mathcal{P}}^{\perp}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\big{|}_{{\mathbb{R}}^{QN}}^{2}$ (5.7)
where ${\rm rank}({\mathcal{P}}^{\perp}[\underline{\theta}_{*}])=QN-K$.
Proof.
Given that ${\rm rank}(DD^{T})=K$ is maximal in $U$, it follows that the matrix $D^{T}D\in{\mathbb{R}}^{K\times K}$ is invertible. Therefore, the expression for the right hand side of (5.2), and (5.4) for the orthoprojector ${\mathcal{P}}$, are well-defined.
We have
$\partial_{s}{\underline{x}}(s)$ $=$ $D[\underline{\theta}(s)]\,\partial_{s}\underline{\theta}(s)$ (5.8)
$=$ $-D[\underline{\theta}(s)](D^{T}[\underline{\theta}(s)]D[\underline{\theta}(s)])^{-1}\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-\underbrace{D[\underline{\theta}(s)](D^{T}[\underline{\theta}(s)]D[\underline{\theta}(s)])^{-1}D^{T}[\underline{\theta}(s)]}_{={\mathcal{P}}[\underline{\theta}(s)]}\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$=$ $-{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]\,,$
where we used (5.2) to pass to the second line. As a consequence,
${\mathcal{P}}^{\perp}[\underline{\theta}(s)]\partial_{s}{\underline{x}}(s)=0\,.$ (5.9)
It follows from the first line in (5.8) that $\partial_{s}\underline{\theta}(s)=0$ implies $\partial_{s}{\underline{x}}(s)=0$.
Let $\underline{\theta}_{*}$ denote a stationary point for (5.2). Then, (5.8) implies that
${\mathcal{P}}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]=0\,.$ (5.10)
Therefore,
${\mathcal{C}}[\underline{\theta}_{*}]$ $=$ $\frac{N}{2}|\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]|^{2}$ (5.11)
$=$ $\frac{N}{2}\Big{(}\big{|}{\mathcal{P}}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\big{|}_{{\mathbb{R}}^{QN}}^{2}+\big{|}{\mathcal{P}}^{\perp}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\big{|}_{{\mathbb{R}}^{QN}}^{2}\Big{)}$
$=$ $\frac{N}{2}\big{|}{\mathcal{P}}^{\perp}[\underline{\theta}_{*}]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}_{*}]]\big{|}_{{\mathbb{R}}^{QN}}^{2}$
as claimed, where ${\rm rank}({\mathcal{P}}^{\perp}[\underline{\theta}_{*}])=QN-K$. ∎
Remark 5.2.
We note that along every orbit $\underline{\theta}(s)$ of (5.2),
$\partial_{s}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]=-\big{|}{\mathcal{P}}[\underline{\theta}(s)]\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\big{|}^{2}\leq 0\,,$ (5.12)
hence the ${\mathcal{L}}^{2}$ cost is monotone decreasing. However, (5.7) is generically strictly positive, and the modified gradient descent does not produce a global minimum of the cost. This matches the results for the standard gradient descent in the underparametrized case obtained in [7].
5.1. Induced gradient flow in output layer
Let $f:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$ denote the map locally represented by ${\underline{x}}:\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]$, so that
$f({\mathbb{R}}^{K})={\mathcal{M}}\subset{\mathbb{R}}^{QN}$ (5.13)
For convenience of presentation, let us here assume that ${\mathcal{M}}$ is an embedded submanifold with constant dimension ${\rm dim}({\mathcal{M}})=K$ everywhere. Then, the tangent subbundle $T{\mathcal{M}}\subset\cup_{p\in{\mathcal{M}}}T_{p}{\mathbb{R}}^{QN}$ is the pushforward vector bundle of $T{\mathbb{R}}^{K}$ under $f$. That is, $T{\mathcal{M}}$ is spanned by sections $f_{*}V\in\Gamma(T{\mathcal{M}})$ for sections $V\in\Gamma(T{\mathbb{R}}^{K})$. In local coordinates, $f_{*}$ is represented by the Jacobian matrix $Df^{\alpha}_{\beta}$ when acting on vector fields, with $\alpha=1,\dots,QN$ and $\beta=1,\dots,K$.
The Euclidean metric on the tangent bundle $T{\mathbb{R}}^{QN}$ induces a Riemannian metric on $T{\mathcal{M}}$, which in turn defines the pullback metric $g$ on $T{\mathbb{R}}^{K}$ given by
$g(V,W)=\sum_{\alpha,\alpha^{\prime},\beta,\beta^{\prime}}\delta_{\alpha,\alpha^{\prime}}Df^{\alpha}_{\beta}Df^{\alpha^{\prime}}_{\beta^{\prime}}V^{\beta}W^{\beta^{\prime}}\;\;,\;\;V,W\in\Gamma({\mathbb{R}}^{K})\,,$ (5.14)
with $\alpha,\alpha^{\prime}=1,\dots,QN$ and $\beta,\beta^{\prime}=1,\dots,K$, in local coordinates where the Euclidean metric has components $\delta_{\alpha,\alpha^{\prime}}$. This is similar as in (2.20). With $[Df^{\alpha}_{\beta}]=D[\underline{\theta}]$, as defined in (1.20), we therefore obtain
$g_{\beta,\beta^{\prime}}[\underline{\theta}]=(D^{T}[\underline{\theta}]D[\underline{\theta}])_{\beta,\beta^{\prime}}\,,$ (5.15)
which endows ${\mathbb{R}}^{K}$ with the pullback Riemannian structure $({\mathbb{R}}^{K},g)$.
The corresponding gradient operator, defined by $d{\mathcal{F}}[V]=g(V,{\rm grad}_{g}{\mathcal{F}})$ for any smooth vector field $V:{\mathbb{R}}^{K}\rightarrow T{\mathbb{R}}^{K}$ and smooth function ${\mathcal{F}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}$ (where $d$ denotes the exterior derivative), has the form
$\sum_{\beta}V^{\beta}\partial_{\beta}{\mathcal{F}}=\sum_{\beta,\beta^{\prime}}g_{\beta,\beta^{\prime}}V^{\beta}({\rm grad}_{g}{\mathcal{F}})^{\beta^{\prime}}$ (5.16)
in local coordinates. Since this holds for all ${\underline{V}}=(V^{1},\dots,V^{K})^{T}\in{\mathbb{R}}^{K}$, we conclude that
${\rm grad}_{g}{\mathcal{F}}$ $=$ $g^{-1}[\underline{\theta}]\nabla_{\underline{\theta}}{\mathcal{F}}$ (5.17)
$=$ $(D^{T}[\underline{\theta}]D[\underline{\theta}])^{-1}\nabla_{\underline{\theta}}{\mathcal{F}}$
in local coordinates $\underline{\theta}\in U_{\alpha}$. We obtain the following result.
Proposition 5.3.
The coordinate invariant expression for the gradient flow generated by $\widetilde{\mathcal{C}}[\underline{\theta}]:={\mathcal{C}}[{\underline{x}}[\underline{\theta}]]$ on $({\mathbb{R}}^{K},g)$ is given by
$\partial_{s}\underline{\theta}(s)=-{\rm grad}_{g}\widetilde{\mathcal{C}}[\underline{\theta}(s)]\,.$ (5.18)
The modified gradient flow (5.2) is the local coordinate representation of (5.18).
In contrast, the standard gradient flow (1.9) is defined with respect to the Euclidean metric in ${\mathbb{R}}^{K}$.
Remark 5.4.
For both the underparametrized case, and the overparametrized case with rank loss, one obtains the constrained Euclidean gradient flow in the output layer of the form (5.3). In the underparametrized case addressed in Theorem 5.1, the projector ${\mathcal{P}}$ is associated to the tangent bundle of the submanifold ${\mathcal{M}}$ in (5.13), and therefore, (5.3) has the structure of a dynamical system with holonomic contraints. In the overparametrized case with rank loss addressed in Theorem 4.1, the projector ${\mathcal{P}}$ is not in general affiliated with a foliation of ${\mathbb{R}}^{QN}$, and therefore, (4.7) has the structure of a dynamical system with non-holonomic constraints.
6. The borderline case $K=QN$
Finally, we remark that in the borderline case $K=QN$ separating the strictly over- and underparametrized situations, we have that $D\in{\mathbb{R}}^{K\times K}$ is a square matrix, and the condition that ${\rm rank}(D)=K$ is maximal implies that it is invertible. Then, the Penrose inverse is equal to the ordinary inverse,
${\rm Pen}[D]=D^{-1}$ (6.1)
and we find that the modified gradient flow (2.1) for the overparametrized system attains the form
$\partial_{s}\underline{\theta}(s)$ $=$ $-D^{-1}[\underline{\theta}(s)]D^{-T}[\underline{\theta}(s)]\nabla_{\underline{\theta}}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\in{\mathbb{R}}^{K}\,.$ (6.2)
On the other hand, the expression for the gradient vector field for the underparametrized system (5.2) can be simplified by use of
$(D^{T}D)^{-1}=D^{-1}D^{-T}\,,$ (6.3)
and thereby reduces to the same expression (6).
Therefore, we conclude that the expressions for the modified gradient flows for the over- and underparametrized systems coincide in the borderline case $K=QN$. In particular, ${\underline{x}}(s)={\underline{x}}[\underline{\theta}(s)]$ then satisfies
$\partial_{s}{\underline{x}}(s)=-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}(s)]\,$ (6.4)
as in the overparametrized case. In particular, this is equivalent to the comparison system given in (1.9), corresponding to the gradient flow in the output layer relative to the Riemannian structure determined by the Euclidean metric.
Acknowledgments: The author thanks Patricia Muñoz Ewald for discussions, the anonymous referees for helpful comments, and gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE.
There are no financial or non-financial competing interests that are directly or indirectly related to this work.